# Subset Balancing and Generalized Subset Sum via Lattices

### 06 Apr 2026 v1

@@ Abstract
We further study the Generalized Subset Sum problem of finding $\mathbf{c}\in C^{n}$ such that $\mathbf{c}\cdot\mathbf{x}=\tau$. For $C=\{-d,\dots,d\}$, we reduce the worst-case problem to a single instance of $\mathrm{CVP}_{\infty}$ in dimension $n+1$. Although no general single exponential time algorithm is known for exact $\mathrm{CVP}_{\infty}$, we show that in the average-case setting, for both $C=\{-d,\dots,d\}$ and $C=\{-d,\dots,d\}\setminus\{0\}$, the embedded instance satisfies a bounded-distance promise with high probability. This yields a deterministic algorithm running in time $\tilde{O}((18\sqrt{2\pi e})^{n})\approx\tilde{O}(2^{6.217n})$.
Our results are complementary to recent representation-based algorithms that outperform meet-in-the-middle for constant-size coefficient sets. Although our exponents are currently larger for small $d$, they do not grow with $|C|$, and they yield clean deterministic guarantees in parameter regimes where lattice methods are more natural than combinatorial enumeration.
@@ 1.1 Generalized Subset Sum, Subset Balancing, and the Meet-in-the-Middle Barrier
This formulation captures, for example, Equal Subset Sum (ESS) when $C=\{-1,0,1\}$, and Partition when $C=\{-1,1\}$ with $\tau=0$. For constant-size coefficient sets $C$, for example $C=[-d:d]:=\{-d,-d+1,\ldots,d\}$ or $C=[\pm d]:=\{-d,\ldots,-1,1,\ldots,d\}$ where $d$ is a fixed integer, Chen et al. [6] obtained running times of the form $|C|^{(1/2-\Omega(1/|C|))n}$ (with high success probability), together with sharp structural results about when solutions exist in the average-case setting. Rewriting this in base 22, the improvement over meet-in-the-middle lies in reducing the constant multiplying $n\log d$ from $1/2$ to $1/2-\Omega(1|C|)$, but the $\log d$ factor in the exponent is preserved. Moreover, the savings $\Omega(1/|C|)$ vanish as $d$ grows, so the improvement becomes negligible for large coefficient sets.
@@
They obtained worst-case algorithms that break the Meet-in-the-Middle barrier for a wide range of constant-size coefficient sets, [d:d][-d:d] for $d\geq 1$ and $\pm d]$ for $d>2$. Their work extends representation-based methods to worst-case inputs via new tools such as coefficient shifting and compatibility certificates, achieving running time $\tilde{O}(|C|^{(1/2-\varepsilon)n})$ for some constant $\varepsilon>0$ depending on $C$. Also, their techniques are are designed for fixed $d$ , as $d$ grows, both the improvement $\varepsilon$ and the algorithmic framework become increasingly unwieldy. For the specific cases computed in [17], the savings $\varepsilon$ decrease from $0.022$ for C=[2:2]C=[-2:2] to $0.005$ for $C=[\pm 3]$, consistent with the interpretation that the exponent remains $\Theta(n\log d)$.
@@ 1.2 Our Contributions
In this work, we first focus on the Subset Balancing Problem with C=[d:d]C=[-d:d] in the worst case. We consider the lattice consisting of all integer solutions $\mathbf{c}$ such that $\mathbf{c}\cdot\mathbf{x}=0$, thereby reducing the task of finding $\mathbf{c}\in C^{n}$ to an instance of $\mathrm{SVP}_{\infty}$ (the Shortest Vector Problem in the $\ell_{\infty}$ norm). However, this lattice is not full rank. To leverage the state-of-the-art deterministic algorithms for $\mathrm{SVP}_{\infty}$ based on the framework of Dadush, Peikert, and Vempala [7] (FOCS 2011), we construct a full-rank lattice by a standard embedding. We further refine the analysis in the $\ell_{\infty}$ setting and obtain the explicit base $6\sqrt{2\pi e}$ appearing in Theorem $1.$ For the randomized setting, we rely on the SVP algorithm of Mukhopadhyay [16], which yields Theorem $1.$
@@
Using an analysis analogous to the one above, we reduce Generalized Subset Sum with C=[d:d]C=[-d:d] to $\mathrm{CVP}_{\infty}$ in the worst case. However, unlike $\mathrm{SVP}_{\infty}$, no general single exponential time algorithm is known for exact $\mathrm{CVP}_{\infty}$ unless the distance between the target vector and the lattice is bounded by a constant multiple of the shortest-vector length [7]. Our second contribution shows that, in the average case regime studied by [6], this bounded distance condition holds with high probability. Specifically, when $\mathbf{x}\sim[0:M-1]^{n}$ and $|\tau|=o(Mn)$, we prove that, with probability $1-e^{-\Omega(n)}$, the distance between the target vector and the lattice is at most 44 times the shortest vector length.
@@
For C=[d:d]C=[-d:d], we obtain the following result (see Theorem˜1.2).
Theorem 1.2 (Algorithms for Average-Case GSS with C=[d:d]C=[-d:d]).
Fix any $d\in\mathbb{N}$ and let C=[d:d]C=[-d:d]. There is a deterministic algorithm for average-case Generalized Subset Sum running in time
For any $M$ and $\tau$ with $|\tau|=o(nM)$, the algorithm succeeds on $(M,\tau,\mathbf{x})$ with probability at least $1-e^{-\Omega(n)}$ (over $\mathbf{x}\sim[0:M-1]^{n}$).
Notably, for $C=[\pm d]$, we have an analogous result, but with a larger failure probability in the following Theorem˜1.3.
Theorem 1.3 (Algorithms for Average-Case GSS with $C=[\pm d]$).
Fix any $d\in\mathbb{N}$ and let $C=[\pm d]$. There is a deterministic algorithm for average-case Generalized Subset Sum running in time
For any $M$ and $\tau$ with $|\tau|=o(nM)$, the algorithm succeeds on $(M,\tau,\mathbf{x})$ with probability at least $1-o_{n}(1)$ (over $\mathbf{x}\sim[0:M-1]^{n}$).
Our results are complementary to the barrier-breaking framework of [6, 17]. Representation-based algorithms excel in the regime where $d$ is a small constant, achieving exponents of the form $(1/2-\varepsilon)\cdot n\log|C|$. In contras, our lattice-based approach operates in a fundamentally different regime: by reducing the problem to lattice problems, we remove the $\log d$ factor from the exponent altogether. This makes our algorithms particularly well-suited to the parameter regime where $d$ is large, and also enables clean deterministic guarantees in settings where randomized hashing is undesirable.
@@ 2 Preliminaries
$\|\mathbf{v}\|_{p}=\begin{cases}\left(\sum_{i=1}^{m}|v_{i}|^{p}\right)^{1/p},&1\leq p<\infty,\\[2.84526pt] \max_{1\leq i\leq m}|v_{i}|,&p=\infty.\end{cases}$ (1)
@@
We assume that all integer inputs ($d,x_{i},M$) are bounded by $2^{n^{O(1)}}$ as in [17]. This is without loss of generality: integers of magnitude $2^{n^{\omega(1)}}$ are already infeasible to read or operate on in polynomial time in any realistic model. Notably, in Section 3 we can suppose $d$ as any integer even a function of $n$.
@@ 2.1 Lattices
Then $\mathcal{L}^{\perp}_{\mathbf{x}}=\ker(\varphi)$, so it is a sublattice of $\mathbb{Z}^{n}$. Since $\varphi$ has rank 11 whenever $\mathbf{x}\neq\mathbf{0}$, it follows that $\ker(\varphi)$ has rank n1n-1.
@@
$\det(\mathcal{L}^{\perp}_{\mathbf{x}})=\frac{\|\mathbf{x}\|_{2}}{\gcd(x_{1},\ldots,x_{n})}.$
@@ 2.2 Convex Geometry
which is the minimum number of translates of $B$ required to cover AA.
@@
$\|\mathbf{x}\|_{K}=\inf\{r\geq 0:\mathbf{x}\in rK\},\quad\mathbf{x}\in\mathbb{R}^{n}.$ (2)
@@
$\langle\mathbf{x},\mathbf{y}\rangle_{\mathbf{A}}=\mathbf{x}\mathbf{A}\mathbf{y}^{t},\quad\|\mathbf{x}\|_{\mathbf{A}}=\sqrt{\mathbf{x}\mathbf{A}\mathbf{x}^{t}}.$ (3)
@@
$E(\mathbf{A},\mathbf{a})=\{\mathbf{x}\in\mathbb{R}^{n}:\|\mathbf{x}-\mathbf{a}\|_{\mathbf{A}}\leq 1\}.$ (4)
@@
$\operatorname{vol}(E(\mathbf{A},\mathbf{a}))=\operatorname{vol}(E(\mathbf{A}))=\operatorname{vol}(B_{2}^{n})\cdot\sqrt{\det(\mathbf{A}^{-1})}.$ (5)
@@ 3.1 Subset Balancing Problems with C=[d:d]C=[-d:d]
Proof of Theorem˜$1.$
Combining Theorems 3.3 and 2.13, there is a deterministic algorithm that solves Subset Balancing with C=[d:d]C=[-d:d] in time $\tilde{O}\!\left(\bigl(6\sqrt{2\pi e}\,\bigr)^{n}\right)\approx\tilde{O}(2^{4.632n})$ for any $d>0$. Moreover, combining Theorem 3.3 with Mukhopadhyay’s $\tilde{O}(2^{2.443n})$ randomized algorithm for $\text{SVP}_{\infty}$ [16], we obtain a randomized algorithm that solves Subset Balancing with C=[d:d]C=[-d:d] in time $\tilde{O}(2^{2.443n})$.∎
@@
Notably, the time complexity of our algorithm is independent with the size of $C$ (ignoring polynomial factor). The ramdomized algorithm with time complexity
@@
We run LLL algorithm with the lattice $\mathcal{L}^{\perp}_{\mathbf{x}}$. By 2.6, we obtain a nonzero vector $\mathbf{v}\in\mathcal{L}^{\perp}_{\mathbf{x}}$ satisfying
@@ 3.2 Generalization of Coefficient Set
We may consider the Minkowski functional induced induced by $K$:
@@
Our algorithm is simply call a SVP oracle (Theorem 2.14) of $\mathcal{L}$ under the norm $\|{\cdot}\|_{K}$. Suppose the SVP oracle returns $\mathbf{v}$, then checks whether $\|\mathbf{v}\|\in K$, if yes, returns $\mathbf{v}$; otherwise, reports that no solution exists.
@@
4 Average Case Algorithms for Generalized Subset Sum

### 24 Apr 2026 v2

@@ Abstract
We further study the Generalized Subset Sum problem of finding $\mathbf{c}\in C^{n}$ such that $\mathbf{c}\cdot\mathbf{x}=\tau$. For $C=\{-d,\dots,d\}$ or $C=\{-d,\dots,d\}\setminus\{0\}$, we reduce the worst-case problem to $\mathrm{CVP}_{\infty}$ in dimension $n+1$. Although no general single-exponential time algorithm is known for exact $\mathrm{CVP}_{\infty}$, we observe that distances in our lattice take only integer values, so an approximate $\mathrm{CVP}_{\infty}$ oracle still suffices. Combining this with the approximate $\mathrm{CVP}_{\infty}$ algorithm yields a deterministic worst-case algorithm running in time $2^{O(n\log\log d)}$, already asymptotically faster than meet-in-the-middle.
In the average-case setting, we demonstrate that for both coefficient sets the embedded $\mathrm{CVP}_{\infty}$ instance satisfies a bounded-distance promise with high probability, which allows us to remove the $\log\log d$ factor altogether. This yields a deterministic algorithm running in time $\tilde{O}((18\sqrt{2\pi e})^{n})\approx\tilde{O}(2^{6.217n})$.
@@ 1.1 Generalized Subset Sum, Subset Balancing, and the Meet-in-the-Middle Barrier
This formulation captures, for example, Equal Subset Sum (ESS) when $C=\{-1,0,1\}$, and Partition when $C=\{-1,1\}$ with $\tau=0$. For constant-size coefficient sets $C$, for example $C=[-d:d]:=\{-d,-d+1,\ldots,d\}$ or $C=[\pm d]:=\{-d,\ldots,-1,1,\ldots,d\}$ where $d$ is a fixed integer, Chen et al. [7] obtained running times of the form $|C|^{(1/2-\Omega(1/|C|))n}$ (with high success probability), together with sharp structural results about when solutions exist in the average-case setting. Rewriting this in base $2$, the improvement over meet-in-the-middle lies in reducing the constant multiplying $n\log d$ from $1/2$ to $1/2-\Omega(1|C|)$, but the $\log d$ factor in the exponent is preserved. Moreover, the savings $\Omega(1/|C|)$ vanish as $d$ grows, so the improvement becomes negligible for large coefficient sets.
@@
They obtained worst-case algorithms that break the Meet-in-the-Middle barrier for a wide range of constant-size coefficient sets, [d:d][-d:d] for $d\geq 1$ and $\pm d]$ for $d>2$. Their work extends representation-based methods to worst-case inputs via new tools such as coefficient shifting and compatibility certificates, achieving running time $\tilde{O}(|C|^{(1/2-\varepsilon)n})$ for some constant $\varepsilon>0$ depending on $C$. Also, their techniques are designed for fixed $d$ , as $d$ grows, both the improvement $\varepsilon$ and the algorithmic framework become increasingly unwieldy. For the specific cases computed in [20], the savings $\varepsilon$ decrease from $0.022$ for C=[2:2]C=[-2:2] to $0.005$ for $C=[\pm 3]$, consistent with the interpretation that the exponent remains $\Theta(n\log d)$.
@@ 1.2 Our Contributions
In this work, we first focus on the Subset Balancing Problem with C=[d:d]C=[-d:d] in the worst case. We consider the lattice consisting of all integer solutions $\mathbf{c}$ such that $\mathbf{c}\cdot\mathbf{x}=0$, thereby reducing the task of finding $\mathbf{c}\in C^{n}$ to an instance of $\mathrm{SVP}_{\infty}$ (the Shortest Vector Problem in the $\ell_{\infty}$ norm). However, this lattice is not full rank. To leverage the state-of-the-art deterministic algorithms for $\mathrm{SVP}_{\infty}$ based on the framework of Dadush, Peikert, and Vempala [9] (FOCS 2011), we construct a full-rank lattice by a standard embedding. We further refine the analysis in the $\ell_{\infty}$ setting and obtain the explicit base $6\sqrt{2\pi e}$ appearing in Theorem 1.1. For the randomized setting, we rely on the SVP algorithm of Mukhopadhyay [19], which yields Theorem 1.1.
@@
Using an analysis analogous to the one above, we reduce Generalized Subset Sum with C=[d:d]C=[-d:d] to $\mathrm{CVP}_{\infty}$ in the worst case. However, unlike $\mathrm{SVP}_{\infty}$, no general single exponential time algorithm is known for exact $\mathrm{CVP}_{\infty}$, and when $C=[\pm d]$ the short vector in the lattice may correspond to an invalid solution in the worst case.
We overcome these by two key observations.
The first observation is that in our lattice, distance from the target to any lattice point can take only integer values, which lets us replace an exact $\mathrm{CVP}_{\infty}$ oracle by an approximate one. Specifically, for the C=[d:d]C=[-d:d] cases, we only need an approximate $\text{CVP}_{\infty}$ oracle with approximation factor $\left(1+\frac{1}{d}-\varepsilon\right)$ for arbitrarily small $\varepsilon>0$. We apply the oracle proposed by Eisenbrand, Hähnle, and Niemeier [11], yielding the overall complexity $2^{O(n\log\log d)}$ in Theorem˜1.2.
When $C=[\pm d]=[-d,-1]\cup[1,d]$, the second observation is that $C^{n}$ can be viewed as a union of $2^{n}$ boxes. For each box, we shift the target vector then call the approximation $\text{CVP}_{\infty}$ oracle. Multiplying by the $2^{n}$ factor yields the same overall time $2^{O(n\log\log d)}$ in Theorem˜1.2.
Theorem 1.2 (Algorithms for Worst-Case GSS with C=[d:d]C=[-d:d] or $\pm d]$).
Given any $d>2,n\in\mathbb{N}$ and let C=[d:d]C=[-d:d] or $\pm d]$, there is a deterministic algorithm for Generalized Subset Sum Problem in running time
$2^{O(n\log\log d)}.$
Finally we study the average cases as [7], in which we remove the $\log\log d$ factor altogether.
We show a bounded distance condition holds with high probability. Specifically, when $\mathbf{x}\sim[0:M-1]^{n}$ and $|\tau|=o(Mn)$, we prove that, with probability $1-e^{-\Omega(n)}$, the distance between the target vector and the lattice is at most $4$ times the shortest vector length. Notably, for $C=[\pm d]$, we have an analogous result, but with a larger failure probability $1-o_{n}(1)$ in the following Theorem˜1.3.
We obtain the following result.
Theorem 1.3 (Algorithms for Average-Case GSS with C=[d:d]C=[-d:d] or $\pm d]$).
Fix $d\in\mathbb{N}$ and let C=[d:d]C=[-d:d] or $\pm d]$. There is a deterministic algorithm for average-case Generalized Subset Sum running in time
For any $M$ and $\tau$ with $|\tau|=o(nM)$, the algorithm succeeds on $(M,\tau,\mathbf{x})$ with probability at least $1-e^{-\Omega(n)}$ when C=[d:d]C=[-d:d] and with probability at least $1-o_{n}(1)$ when $C=[\pm d]$ (over $\mathbf{x}\sim[0:M-1]^{n}$).
Table 1: Comparison of algorithms for Subset Balancing (SBP) and Generalized Subset Sum (GSS) with C=[d:d]C=[-d:d]. The column “Exp. in dd” summarizes how the exponent of the running time scales with $d$. Here $\tilde{O}$ suppresses $\operatorname{poly}(n)$ factors. Analogous results for $C=[\pm d]$ are stated in the text.
Problem Reference Setting Type Running Time Exp. in $d$
SBP MitM [13] worst det $\tilde{O}\!\bigl(|C|^{n/2}\bigr)$ $\Theta(n\log d)$
[20] worst rand $\tilde{O}\!\bigl(|C|^{(1/2-\varepsilon)n}\bigr)^{\dagger}$ $\Theta(n\log d)$
Ours worst det $\tilde{O}(2^{4.632n})$ $O(n)$
Ours worst rand $\tilde{O}(2^{2.443n})$ $O(n)$
Ours worst det $\operatorname{poly}(n,\log d)^{\ddagger}$
GSS MitM [13] worst det $\tilde{O}\!\bigl(|C|^{n/2}\bigr)$ $\Theta(n\log d)$
[7] avg rand $\tilde{O}\!\bigl(|C|^{(1/2-\varepsilon)n}\bigr)^{\dagger}$ $\Theta(n\log d)$
Ours worst det $2^{O(n\log\log d)}$ $O(n\log\log d)$
Ours avg det $\tilde{O}(2^{6.217n})$ $O(n)$
$\varepsilon=\varepsilon(|C|)>0$ depends on $|C|$ and satisfies $\varepsilon\to 0$ as $d\to\infty$.
 Requires $d$ to be sufficiently large; see Theorem 3.5.
@@
Our results are complementary to the barrier-breaking framework of [7, 20]. Representation-based algorithms excel in the regime where $d$ is a small constant, achieving exponents of the form $(1/2-\varepsilon)\cdot n\log|C|$. In contrast, our lattice-based approach operates in a fundamentally different regime: by reducing the problem to lattice problems, we remove the $\log d$ factor from the exponent altogether. This makes our algorithms particularly well-suited to the parameter regime where $d$ is large, and also enables clean deterministic guarantees in settings where randomized hashing is undesirable.
@@ 2 Preliminaries
$\|\mathbf{v}\|_{p}=\begin{cases}\left(\sum_{i=1}^{m}|v_{i}|^{p}\right)^{1/p},&1\leq p<\infty,\\[2.84526pt] \max_{1\leq i\leq m}|v_{i}|,&p=\infty.\end{cases}$
@@
We assume that all integer inputs ($d,x_{i},M$) are bounded by $2^{n^{O(1)}}$ as in [20]. This is without loss of generality: integers of magnitude $2^{n^{\omega(1)}}$ are already infeasible to read or operate on in polynomial time in any realistic model. Notably, in Sections˜3 and 4 we can suppose $d$ as any integer even a function of $n$.
@@ 2.1 Lattices
Definition 2.8 ($\gamma$-Approximate Closest Vector Problem ($\gamma$-CVPp)).
Let $p\in[1,\infty],\gamma>1$. Given a lattice $\mathcal{L}\subset\mathbb{R}^{m}$ and a target vector $\mathbf{t}\in\mathbb{R}^{m}$, the $\gamma$-Approximate Closest Vector Problem in the $\ell_{p}$ norm ($\gamma$-CVPp) asks to find $\mathbf{v}\in\mathcal{L}$ satisfying
$\|\mathbf{v}-\mathbf{t}\|_{p}\leq\gamma\min_{\mathbf{w}\in\mathcal{L}}\|\mathbf{t}-\mathbf{w}\|_{p}.$
@@
Then $\mathcal{L}^{\perp}_{\mathbf{x}}=\ker(\varphi)$, so it is a sublattice of $\mathbb{Z}^{n}$. Since $\varphi$ has rank $1$ whenever $\mathbf{x}\neq\mathbf{0}$, it follows that $\ker(\varphi)$ has rank n1n-1.
@@
$\det(\mathcal{L}^{\perp}_{\mathbf{x}})=\frac{\|\mathbf{x}\|_{2}}{\gcd(x_{1},\ldots,x_{n})}.\qed$
@@ 2.2 Convex Geometry
which is the minimum number of translates of $B$ required to cover $A$.
@@
$\|\mathbf{x}\|_{K}=\inf\{r\geq 0:\mathbf{x}\in rK\},\quad\mathbf{x}\in\mathbb{R}^{n}.$
@@
$\langle\mathbf{x},\mathbf{y}\rangle_{\mathbf{A}}=\mathbf{x}\mathbf{A}\mathbf{y}^{t},\quad\|\mathbf{x}\|_{\mathbf{A}}=\sqrt{\mathbf{x}\mathbf{A}\mathbf{x}^{t}}.$
@@
$E(\mathbf{A},\mathbf{a})=\{\mathbf{x}\in\mathbb{R}^{n}:\|\mathbf{x}-\mathbf{a}\|_{\mathbf{A}}\leq 1\}.$
@@
$\operatorname{vol}(E(\mathbf{A},\mathbf{a}))=\operatorname{vol}(E(\mathbf{A}))=\operatorname{vol}(B_{2}^{n})\cdot\sqrt{\det(\mathbf{A}^{-1})}.$
@@ 12 $d\leftarrow(1+\varepsilon)\,d$;
$\tilde{O}\!\left(\bigl(6\sqrt{2\pi e}\,\bigr)^{n}\right)\approx\tilde{O}(2^{4.632n}).\qed$
@@ 3.1 Subset Balancing Problems with C=[d:d]C=[-d:d]
Proof of Theorem˜1.1.
Combining Theorems 3.3 and 2.14, there is a deterministic algorithm that solves Subset Balancing with C=[d:d]C=[-d:d] in time $\tilde{O}\!\left(\bigl(6\sqrt{2\pi e}\,\bigr)^{n}\right)\approx\tilde{O}(2^{4.632n})$ for any $d>0$. Moreover, combining Theorem 3.3 with Mukhopadhyay’s $\tilde{O}(2^{2.443n})$ randomized algorithm for $\text{SVP}_{\infty}$ [19], we obtain a randomized algorithm that solves Subset Balancing with C=[d:d]C=[-d:d] in time $\tilde{O}(2^{2.443n})$.∎
@@
Notably, the time complexity of our algorithm is independent with the size of $C$ (ignoring polynomial factor). The randomized algorithm with time complexity
@@
We run LLL algorithm with the lattice $\mathcal{L}^{\perp}_{\mathbf{x}}$. By Lemma˜2.6, we obtain a nonzero vector $\mathbf{v}\in\mathcal{L}^{\perp}_{\mathbf{x}}$ satisfying
@@ 3.2 Generalization of Coefficient Set
We may consider the Minkowski functional induced by $K$:
@@
Our algorithm simply calls an SVP oracle (Theorem 2.15) of $\mathcal{L}$ under the norm $\|{\cdot}\|_{K}$. Suppose the SVP oracle returns $\mathbf{v}$, then checks whether $\|\mathbf{v}\|\in K$, if yes, returns $\mathbf{v}$; otherwise, reports that no solution exists.
@@
4 Worst Case Algorithms for Generalized Subset Sum